# FaceShield: Defending Facial Image against Deepfake Threats

### v2

@@ 1 Introduction
Figure 1: Protecting Face during Deepfake using FaceShield. Pure images are vulnerable to face swapping, allowing the target image’s face to be easily reflected. In contrast, images protected by FaceShield conceal facial feature from deepfake.
@@
The advancement of deepfake technology and improved accessibility [28, 36, 71, 4, 55] has led to significant transformations in modern society. Due to the ease of face swapping, it has been applied across various fields, providing both entertainment and convenience. However, its powerful capability to generate realistic content has also enabled malicious users to exploit it for criminal purposes, leading to the creation of fake news and various societal problems.
To address the growing concerns surrounding deepfake technology, various countermeasures have been explored, which can be broadly divided into two categories. The first is deepfake detection techniques [46, 18, 16, 54, 6], which act as passive defenses by classifying whether content is synthetic or authentic. While effective for authenticity verification, these offer only binary results and fail to address advanced threats, such as crimes using realistic fakes. In contrast, proactive defense strategies offer a more comprehensive solution. These approaches involve embedding imperceptible adversarial perturbation into face images to prevent the protected face from being effectively processed by deepfakes. However, most previous research [62, 40, 10, 19, 53] has concentrated on GAN-based models, often targeting individual models, which limits effectiveness against emerging DM-based deepfakes [25, 68, 50, 61]. Although significant research exists on image protection within DMs [32, 43, 42, 31, 60, 7] for image editing, the focus has primarily been on attacking the noising and denoising processes when an image is used as a query (Fig.2a). This leads to targeting the encoder or predicted noise post-UNet processing. However, we observe that such strategies are ineffective for DM-based deepfake models, where the source image influences the output in the form of key-value pairs through attention mechanisms (Fig.2b).
@@
For DM-based deepfake attacks, we leverage the structural properties in which the conditioning image is embedded and integrated into the denoising UNet through attention mechanisms. By utilizing the IP-Adapter [61], commonly employed for inpainting, we extract effective adversarial noise from the embedding of the conditioning image. This perturbation effectively disrupts the propagation of the conditioning process, ensuring that the final output does not replicate the features of the protected image.
To enhance the generalizability of our approach, we target two commonly used facial feature extractors. First, we attack the MTCNN model [67], which uses a cascade pyramid architecture to achieve superior performance and robust detection capabilities. Due to this, it is widely adopted not only in deepfake generation but also across various applications. We leverage the fact that the model scales images to different sizes during face detection. Our perturbation is designed to ensure robustness across various scaling factors and interpolation modes (e.g., $\mathtt{BILINEAR}$, $\mathtt{AREA}$), leading to superior performance compared to existing methods [24, 65]. Additionally, we target ArcFace [9], a widely adopted pre-trained model for facial feature extraction in deepfake applications. By incorporating both methods into our work, we ensure that our approach disrupts a range of deepfakes commonly used for facial landmark detection and feature extraction, thereby improving the overall robustness of our method against various deepfake systems.
@@ 2 Related Work
Deepfake techniques. With advancements in generative models, deepfake technology has evolved into a specialized field focused on facial synthesis. Previous deepfake models, primarily based on GANs, generally follow a three-stage process: face detection and localization, feature extraction, and face swapping. Among these, studies such as [71, 58, 14, 59, 64] employ MTCNN [67] for face detection and landmark extraction, while the majority of deepfake models, including [28, 4, 71, 27, 58, 72], leverage ArcFace [9] for identity feature extraction. These steps are similarly employed in DM-based deepfakes that have emerged with the progress of diffusion models. Notable examples, including [25, 68, 50], integrate [9] to maintain identity consistency. However, recent work has focused on leveraging the capabilities of diffusion models to develop face-swapping methods [61] that achieve high performance without explicitly following previous approaches.
Enhancing model transferability. In the research on adversarial attacks, various attempts have been made to improve transferability. [11, 56] proposed the model ensemble technique, generating adversarial examples using multiple models to enhance their effectiveness on unseen models. [57] introduced a method that selectively utilizes specific layers within a model to improve transferability. Similarly, [21, 20, 69] investigated techniques that manipulate intermediate layer feature distributions or amplify activation values to prevent adversarial noise from overfitting to a particular model. Furthermore, [3, 66] explored the use of multiple pre-trained backbones within similar model architectures to enhance transferability across different backbone networks.
@@ 3.2 Conditioned Face Attack
We now describe our approach to protecting images, specifically by disrupting the effective transfer of information when they are used as conditioning inputs in DMs. The core of our approach is to effectively interfere with key information using minimal noise, while also ensuring that the model does not overfit by accessing only a minimal number of layers to obtain gradients. To achieve this, we propose two methods that target both the initial projection phase and the final attention mechanism during the image conditioning process within latent diffusion models [39].
Face projector attack. When images are used as conditioning inputs, they are firstly transformed into an embedding vector through a pre-trained model [38]. In this method, we access only the topmost layer $\mathcal{P}$ of the model to disrupt the projection process, causing the image to be projected with incorrect information at the initial stage. For the attack loss function, we consider that converging to a single target value might not ensure consistent convergence speeds or balanced performance. Given that one of our main goals is to design noise applicable to various images, we design our approach to induce random divergence based on the input image, using the $\mathcal{L}_{1}$ loss function in this process:
@@
Attention disruption attack. We also focus on identifying the core vectors within the denoising UNet that are most sensitive to conditional inputs. Initially, we analyze the influence of cross-attention across each UNet layer. Based on prior research [49], which shows that different cross-attention layers respond variably to conditioning information, we investigate the impact on perturbation performance for each region. Our findings lead to the conclusion that targeting attacks near mid-layers produces more significant disruption in qualitative metrics compared to using only the up-down layers or the entire layers, as supported by our experimental results in Fig.6. Based on these insights, we propose a novel approach that specifically targets mid-layers during the attack on the diffusion process.
@@
$A_{\mathtt{var}}=\frac{1}{\mathtt{seq}}\sum_{i=1}^{\mathtt{seq}}\left(A_{\mathtt{map}}[:,i,:]-\bar{A}_{\mathtt{map}}\right)^{2}\in\mathbb{R}^{h\times\mathtt{res}},$ (5)
where $\bar{A}_{\mathtt{map}}=\frac{1}{\mathtt{seq}}\sum_{i^{\prime}=1}^{\mathtt{seq}}A_{\mathtt{map}}[:,i^{\prime},:]$ is the mean of attention map across the $\mathtt{seq}$ dimension.
@@ 3.3 Facial Feature Extractor Attack
Identity attack. To effectively disrupt the accurate reflection of source face information, we target the ArcFace $\mathcal{A}$ models [9], which are face identity embedding models widely used in deepfake systems. To improve transferability, we ensemble the most commonly used pre-trained backbones within these models. Since $\mathcal{A}$ represents feature vectors extracted from the same person’s face as vectors pointing in similar directions, we designed our approach to induce divergence from the original image $x$ by employing cosine similarity loss, thereby effectively obscuring the relevant identity information:
@@
$\mathcal{L}_{\mathtt{total}}=$ $\lambda_{\mathtt{proj}}\mathcal{L}_{\mathtt{proj}}+\lambda_{\mathtt{attn}}\mathcal{L}_{\mathtt{attn}}$
@@
Figure 4: Qualitative Results. Protection performance across various deepfake models when our adversarial noise is applied. Models [50, 61] highlighted in the orange box typically exhibit facial distortions due to the influence described in Sec. 3.2, while those [68, 25, 4, 14] in the blue box display newly generated faces that diverge from the source image, attributed to the impact detailed in Sec. 3.3.
@@ 3.4 Enhanced Noise Update
Input: image $x$, steps $N$, noise clamp $\epsilon$, step size $\alpha$, MTCNN detection threshold $t_{\mathtt{prob}}$, threshold weight $\beta$, CLIP Image Projector $\mathcal{P}$, Mid-layer cross-attention variance in Stable Diffusion $A_{\mathtt{var}}^{\prime}$, MTCNN P-Network $\mathcal{T}$, ArcFace $\mathcal{A}$
@@ 12         $x_{\mathtt{adv}}\leftarrow x+\text{clip}(x_{\mathtt{adv}}-x,-\epsilon,\epsilon)$
Gaussian blur. To enhance noise imperceptibility, we introduce a technique that constrains variations between adjacent regions, addressing the limitations of PGD methods (see Eq. (3)) that only regulate overall noise magnitude. This stems from the observation that differences between neighboring pixels can be as perceptible as the total noise itself. To achieve this, we utilize the Sobel operator [22] to emphasize areas of rapid intensity change, generating a mask $M_{\mathtt{sob}}$ that highlights image boundaries. Gaussian blur $\mathcal{G}(\cdot)$ is then applied selectively to these regions during noise updates, ensuring smoother transitions between adjacent pixels and maintaining a consistent visual appearance:
@@
Low pass filtering. To minimize information loss when saving images in $\mathtt{JPEG}$ format and ensure robustness to bit reduction during the compression process [13], low-frequency components are utilized. At each iteration, the newly updated adversarial noise $\delta_{\mathtt{rgb}}\in\mathbb{R}^{c\times h\times w}$, where $c$ (channel), $h$ (height), and $w$ (width), undergoes a padding operation and patchification according to a predefined patch size $p$. Then a DCT transform [1] is applied to each patch and channel, resulting in $\delta_{\mathtt{dct}}\in\mathbb{R}^{c\times h^{\prime}\times w^{\prime}\times p\times p}$, where $h^{\prime}=h/p$ and $w^{\prime}=w/p$, in the frequency domain. Using a low-pass filtering mask $M_{\mathtt{lp}}$, only the low-frequency components of the noise are retained. The noise $\delta^{\prime}_{\mathtt{rgb}}\in\mathbb{R}^{c\times h\times w}$ is then reconstructed back into the RGB domain through an inverse transformation. The effectiveness of this approach is demonstrated through the experimental results presented in the supplementary material, and the overall operation of FaceShield is described in Algorithm 1.
[…]
Table 1: Comparison of perturbation effectiveness among baseline methods on four deepfake models using the CelebA-HQ [23] and VGGFace2-HQ [5] datasets. Our method exhibits the largest distortion in image quality (L2, PSNR) and source similarity (ISM), as well as in human evaluation (HE). Results on JPEG-compressed images (Quality factor 75) further confirm robust protection under compression.
@@ 4.1 Setups
Evaluation details. For a fair performance comparison, we use open-source baseline [32, 31, 42, 60] and apply noise to the same dataset under identical hyperparameter settings. The corresponding results are presented in Table 2, while Table 1 provides a performance comparison on diffusion-based deepfakes [25, 68, 50, 61]. The extensibility experiments on GAN-based models [4, 14] are shown in Table 4, where, in the absence of existing attack methods for these models, we validate FaceShield’s effectiveness through comparisons with the original images. In cases where the feature extractor fails to detect a face, we adjust the generation process to exclude facial features during reconstruction. Detailed descriptions and an analysis of the resources are provided in the supplementary material.
Datasets. We evaluate our method using two datasets: CelebA-HQ [23] and VGGFace2-HQ [5], both of which have been used in previous studies [4, 50, 14]. The former is the high-resolution version of CelebA, containing 30,000 celebrity face images, while the latter is the high-resolution version of VGGFace2, consisting of 3.3 million face images from 9,131 unique identities. For our experiments, we randomly select 200 identities from each dataset, using 100 images for the source and 100 images for the target.

### v3

@@ 1 Introduction
Figure 1: Protecting Face during Deepfake using FaceShield. Pure images are vulnerable to face swapping, allowing the target image’s face to be easily reflected. In contrast, images protected by FaceShield conceal facial feature from deepfake. Code is available here: https://github.com/kuai-lab/iccv25_faceshield
@@
The advancement of deepfake technology and improved accessibility [li2019faceshifter, Liu2020DeepfacelabIF, zhu2021one, chen2020simswap, wang2021hififace] has led to significant transformations in modern society. Due to the ease of face swapping, it has been applied across various fields, providing both entertainment and convenience. However, its powerful capability to generate realistic content has also enabled malicious users to exploit it for criminal purposes, leading to the creation of fake news and various societal problems.
To address the growing concerns surrounding deepfake technology, various countermeasures have been explored, which can be broadly divided into two categories. The first is deepfake detection techniques [tan2023learning, huang2023implicit, guo2023hierarchical, wang2023noise, choi2024exploiting], which act as passive defenses by classifying whether content is synthetic or authentic. While effective for authenticity verification, these offer only binary results and fail to address advanced threats, such as crimes using realistic fakes. In contrast, proactive defense strategies offer a more comprehensive solution. These approaches involve embedding imperceptible adversarial perturbation into face images to prevent the protected face from being effectively processed by deepfakes. However, most previous research [yeh2020disrupting, ruiz2020disrupting, dong2021visually, huang2022cmua, wang2022anti] has concentrated on GAN-based models, often targeting individual models, which limits effectiveness against emerging DM-based deepfakes [kim2212diffface, zhao2023diffswap, wang2024face, ye2023ip]. Although significant research exists on image protection within DMs [liang2023adversarial, shan2023glaze, salman2023raising, liang2023mist, xue2023toward, choi2024diffusionguard] for image editing, the focus has primarily been on attacking the noising and denoising processes when an image is used as a query (Fig.2a). This leads to targeting the encoder or predicted noise post-UNet processing. However, we observe that such strategies are ineffective for DM-based deepfake models, where the source image influences the output in the form of key-value pairs through attention mechanisms (Fig.2b).
@@
For DM-based deepfake attacks, we leverage the structural properties in which the conditioning image is embedded and integrated into the denoising UNet through attention mechanisms. By utilizing the IP-Adapter [ye2023ip], commonly employed for inpainting, we extract effective adversarial noise from the embedding of the conditioning image. This perturbation effectively disrupts the propagation of the conditioning process, ensuring that the final output does not replicate the features of the protected image.
To enhance the generalizability of our approach, we target two commonly used facial feature extractors. First, we attack the MTCNN model [zhang2016joint], which uses a cascade pyramid architecture to achieve superior performance and robust detection capabilities. Due to this, it is widely adopted not only in deepfake generation but also across various applications. We leverage the fact that the model scales images to different sizes during face detection. Our perturbation is designed to ensure robustness across various scaling factors and interpolation modes (e.g., $\mathtt{BILINEAR}$, $\mathtt{AREA}$), leading to superior performance compared to existing methods [kaziakhmedov2019real, zhang2022multi]. Additionally, we target ArcFace [deng2019arcface], a widely adopted pre-trained model for facial feature extraction in deepfake applications. By incorporating both methods into our work, we ensure that our approach disrupts a range of deepfakes commonly used for facial landmark detection and feature extraction, thereby improving the overall robustness of our method against various deepfake systems.
@@ 2 Related Work
Deepfake techniques. With advancements in generative models, deepfake technology has evolved into a specialized field focused on facial synthesis. Previous deepfake models, primarily based on GANs, generally follow a three-stage process: face detection and localization, feature extraction, and face swapping. Among these, studies such as [zhu2021one, xu2022region, gao2021information, xu2022high, yuan2023reliableswap] employ MTCNN [zhang2016joint] for face detection and landmark extraction, while the majority of deepfake models, including [li2019faceshifter, chen2020simswap, zhu2021one, li2021faceinpainter, xu2022region, zhu2024stableswap], leverage ArcFace [deng2019arcface] for identity feature extraction. These steps are similarly employed in DM-based deepfakes that have emerged with the progress of diffusion models. Notable examples, including [kim2212diffface, zhao2023diffswap, wang2024face], integrate [deng2019arcface] to maintain identity consistency. However, recent work has focused on leveraging the capabilities of diffusion models to develop face-swapping methods [ye2023ip] that achieve high performance without explicitly following previous approaches.
Enhancing model transferability. In the research on adversarial attacks, various attempts have been made to improve transferability. [dong2018boosting, wang2021feature] proposed the model ensemble technique, generating adversarial examples using multiple models to enhance their effectiveness on unseen models. [xie2019improving] introduced a method that selectively utilizes specific layers within a model to improve transferability. Similarly, [inkawhich2020transferable, huang2019enhancing, zhou2018transferable] investigated techniques that manipulate intermediate layer feature distributions or amplify activation values to prevent adversarial noise from overfitting to a particular model. Furthermore, [chen2024rethinking, zhang2024on] explored the use of multiple pre-trained backbones within similar model architectures to enhance transferability across different backbone networks.
@@ 3.2 Conditioned Face Attack
We now describe our approach to protecting images, specifically by disrupting the effective transfer of information when they are used as conditioning inputs in DMs. The core of our approach is to effectively interfere with key information using minimal noise, while also ensuring that the model does not overfit by accessing only a minimal number of layers to obtain gradients. To achieve this, we propose two methods that target both the initial projection phase and the final attention mechanism during the image conditioning process within latent diffusion models [Rombach_2022_CVPR].
Face projector attack. When images are used as conditioning inputs, they are firstly transformed into an embedding vector through a pre-trained model [radford2021learning]. In this method, we access only the topmost layer $\mathcal{P}$ of the model to disrupt the projection process, causing the image to be projected with incorrect information at the initial stage. For the attack loss function, we consider that converging to a single target value might not ensure consistent convergence speeds or balanced performance. Given that one of our main goals is to design noise applicable to various images, we design our approach to induce random divergence based on the input image, using the $\mathcal{L}_{1}$ loss function in this process:
@@
Attention disruption attack. We also focus on identifying the core vectors within the denoising UNet that are most sensitive to conditional inputs. Initially, we analyze the influence of cross-attention across each UNet layer. Based on prior research [voynov2023p+], which shows that different cross-attention layers respond variably to conditioning information, we investigate the impact on perturbation performance for each region. Our findings lead to the conclusion that targeting attacks near mid-layers produces more significant disruption in qualitative metrics compared to using only the up-down layers or the entire layers, as supported by our experimental results in Fig.6. Based on these insights, we propose a novel approach that specifically targets mid-layers during the attack on the diffusion process.
@@
$A_{\mathtt{var}}=\frac{1}{\mathtt{seq}}\sum_{i=1}^{\mathtt{seq}}\left(A_{\mathtt{map}}[:,:,i]-\bar{A}_{\mathtt{map}}\right)^{2}\in\mathbb{R}^{h\times\mathtt{res}},$ (5)
where $\bar{A}_{\mathtt{map}}=\frac{1}{\mathtt{seq}}\sum_{i^{\prime}=1}^{\mathtt{seq}}A_{\mathtt{map}}[:,:,i^{\prime}]$ is the mean of attention map across the $\mathtt{seq}$ dimension.
@@ 3.3 Facial Feature Extractor Attack
Identity attack. To effectively disrupt the accurate reflection of source face information, we target the ArcFace $\mathcal{A}$ models [deng2019arcface], which are face identity embedding models widely used in deepfake systems. To improve transferability, we ensemble the most commonly used pre-trained backbones within these models. Since $\mathcal{A}$ represents feature vectors extracted from the same person’s face as vectors pointing in similar directions, we designed our approach to induce divergence from the original image $x$ by employing cosine similarity loss, thereby effectively obscuring the relevant identity information:
@@
$\mathcal{L}_{\mathtt{total}}=\$$\lambda_{\mathtt{proj}}\mathcal{L}_{\mathtt{proj}}+\lambda_{\mathtt{attn}}\mathcal{L}_{\mathtt{attn}}$
@@
Figure 4: Qualitative Results. Protection performance across various deepfake models when our adversarial noise is applied. Models [wang2024face, ye2023ip] highlighted in the orange box typically exhibit facial distortions due to the influence described in Sec. 3.2, while those [zhao2023diffswap, kim2212diffface, chen2020simswap, gao2021information] in the blue box display newly generated faces that diverge from the source image, attributed to the impact detailed in Sec. $3.$
@@ 3.4 Enhanced Noise Update
Input: image $x$, steps NN, noise clamp $\epsilon$, step size $\alpha$, MTCNN detection threshold $t_{\mathtt{prob}}$, threshold weight $\beta$, CLIP Image Projector $\mathcal{P}$, Mid-layer cross-attention variance in Stable Diffusion $A_{\mathtt{var}}^{\prime}$, MTCNN P-Network $\mathcal{T}$, ArcFace $\mathcal{A}$
@@ 12     $x_{\mathtt{adv}}\leftarrow x+\text{clip}(x_{\mathtt{adv}}-x,-\epsilon,\epsilon)$
Gaussian blur. To enhance noise imperceptibility, we introduce a technique that constrains variations between adjacent regions, addressing the limitations of PGD methods (see Eq. (3)) that only regulate overall noise magnitude. This stems from the observation that differences between neighboring pixels can be as perceptible as the total noise itself. To achieve this, we utilize the Sobel operator [kanopoulos1988design] to emphasize areas of rapid intensity change, generating a mask $M_{\mathtt{sob}}$ that highlights image boundaries. Gaussian blur $\mathcal{G}(\cdot)$ is then applied selectively to these regions during noise updates, ensuring smoother transitions between adjacent pixels and maintaining a consistent visual appearance:
@@
Low pass filtering. To minimize information loss when saving images in $\mathtt{JPEG}$ format and ensure robustness to bit reduction during the compression process [dziugaite2016study], low-frequency components are utilized. At each iteration, the newly updated adversarial noise $\delta_{\mathtt{rgb}}\in\mathbb{R}^{c\times h\times w}$, where $c$ (channel), $h$ (height), and $w$ (width), undergoes a padding operation and patchification according to a predefined patch size $p$. Then a DCT transform [1672377] is applied to each patch and channel, resulting in $\delta_{\mathtt{dct}}\in\mathbb{R}^{c\times h^{\prime}\times w^{\prime}\times p\times p}$, where $h^{\prime}=h/p$ and $w^{\prime}=w/p$, in the frequency domain. Using a low-pass filtering mask $M_{\mathtt{lp}}$, only the low-frequency components of the noise are retained. The noise $\delta^{\prime}_{\mathtt{rgb}}\in\mathbb{R}^{c\times h\times w}$ is then reconstructed back into the RGB domain through an inverse transformation. The effectiveness of this approach is demonstrated through the experimental results presented in the supplementary material, and the overall operation of FaceShield is described in Algorithm 1.
  Model DiffFace [kim2212diffface] DiffSwap [zhao2023diffswap] FaceSwap [wang2024face] IP-Adapter [ye2023ip] Dataset CelebA-HQ [karras2017progressive] Method $L_{2}$ $\uparrow$ ISM $\downarrow$ PSNR $\downarrow$ HE $\uparrow$ $L_{2}$ $\uparrow$ ISM $\downarrow$ PSNR $\downarrow$ HE $\uparrow$ $L_{2}$ $\uparrow$ ISM $\downarrow$ PSNR $\downarrow$ HE $\uparrow$ $L_{2}$ $\uparrow$ ISM $\downarrow$ PSNR $\downarrow$ HE $\uparrow$ AdvDM [liang2023adversarial] 0.021 0.471 39.368 4.22 0.068 0.199 28.362 4.68 0.303 0.245 21.615 4.52 0.207 0.235 25.332 2.76 Mist [liang2023mist] 0.021 0.468 39.443 3.94 0.067 0.201 28.384 4.18 0.287 0.230 22.263 4.78 0.152 0.265 28.213 4.26 PhotoGuard [salman2023raising] 0.022 0.469 39.194 3.82 0.068 0.201 28.292 4.58 0.282 0.238 22.316 4.44 0.153 0.268 28.101 4.44 SDST [xue2023toward] 0.021 0.470 39.512 4.08 0.067 0.207 28.383 5.04 0.274 0.261 22.582 4.68 0.147 0.273 28.440 4.32 Ours 0.044 0.243 32.052 5.76 0.072 0.163 27.833 6.20 0.336 0.194 20.759 6.16 0.350 0.072 20.266 6.60 Ours (Q=75) 0.043 0.259 32.259 - 0.070 0.169 28.034 - 0.317 0.209 21.286 - 0.326 0.112 20.867 - Dataset VGGFace2-HQ [simswapplusplus] Method $L_{2}$ $\uparrow$ ISM $\downarrow$ PSNR $\downarrow$ HE $\uparrow$ $L_{2}$ $\uparrow$ ISM $\downarrow$ PSNR $\downarrow$ HE $\uparrow$ $L_{2}$ $\uparrow$ ISM $\downarrow$ PSNR $\downarrow$ HE $\uparrow$ $L_{2}$ $\uparrow$ ISM $\downarrow$ PSNR $\downarrow$ HE $\uparrow$ AdvDM [liang2023adversarial] 0.042 0.479 33.064 3.68 0.105 0.215 24.769 4.78 0.419 0.361 18.596 4.38 0.251 0.271 23.250 2.36 Mist [liang2023mist] 0.041 0.478 33.215 4.26 0.102 0.227 24.964 3.94 0.379 0.259 19.626 4.50 0.181 0.291 26.070 4.10 PhotoGuard [salman2023raising] 0.043 0.479 32.938 3.96 0.110 0.215 24.272 4.18 0.373 0.266 19.655 4.14 0.180 0.294 26.157 3.82 SDST [xue2023toward] 0.041 0.483 33.242 5.30 0.107 0.225 24.506 4.58 0.359 0.258 19.996 4.14 0.166 0.292 26.784 4.06 Ours 0.062 0.278 29.204 6.10 0.113 0.177 24.054 6.12 0.453 0.237 17.919 6.16 0.382 0.112 19.478 6.42 Ours (Q=75) 0.060 0.308 29.435 - 0.112 0.185 24.201 - 0.421 0.237 18.573 - 0.377 0.167 19.618 -
Table 1: Comparison of perturbation effectiveness among baseline methods on four deepfake models using the CelebA-HQ [karras2017progressive] and VGGFace2-HQ [simswapplusplus] datasets. Our method exhibits the largest distortion in image quality (L2, PSNR) and source similarity (ISM), as well as in human evaluation (HE). Results on JPEG-compressed images (Quality factor 75) further confirm robust protection under compression.
@@ 4.1 Setups
Evaluation details. For a fair performance comparison, we use open-source baseline [liang2023adversarial, liang2023mist, salman2023raising, xue2023toward] and apply noise to the same dataset under identical hyperparameter settings. The corresponding results are presented in Table 2, while Table 1 provides a performance comparison on diffusion-based deepfakes [kim2212diffface, zhao2023diffswap, wang2024face, ye2023ip]. The extensibility experiments on GAN-based models [chen2020simswap, gao2021information] are shown in Table 4, where, in the absence of existing attack methods for these models, we validate FaceShield’s effectiveness through comparisons with the original images. In cases where the feature extractor fails to detect a face, we adjust the generation process to exclude facial features during reconstruction. Detailed descriptions and an analysis of the resources are provided in the supplementary material.
Datasets. We evaluate our method using two datasets: CelebA-HQ [karras2017progressive] and VGGFace2-HQ [simswapplusplus], both of which have been used in previous studies [chen2020simswap, wang2024face, gao2021information]. The former is the high-resolution version of CelebA, containing 30,000 celebrity face images, while the latter is the high-resolution version of VGGFace2, consisting of 3.3 million face images from 9,131 unique identities. For our experiments, we randomly select 200 identities from each dataset, using 100 images for the source and 100 images for the target.